OpenStudy (korosh23):

Solve algebraically for the value of x in the interval 0 (smaller or equal to) x (less than) 2pi

OpenStudy (korosh23):

\[4sinxcosx = 2-2\cos^2x\]

OpenStudy (korosh23):

@Hero

OpenStudy (mathmale):

Why not simplify matters by dividing every term by 2? Not strictly necessary, but nice.

OpenStudy (mathmale):

Next, look at the right side. Can you think of an identity for it?

OpenStudy (korosh23):

yes sin^2x

OpenStudy (mathmale):

that might be a possibility. Were it not for that coefficient 2 on the left side, I'd try rewriting the right side in the form (sin x)^2. Your turn. Ideas?

OpenStudy (korosh23):

so you mean divide evertything by 2 first?

OpenStudy (mathmale):

Is there an identity for 2 sin x cos x?

OpenStudy (mathmale):

I did suggest you divide every term of the original equation by 2, yes. type out what you have now as a result.

OpenStudy (korosh23):

ok a moment

OpenStudy (korosh23):

2 sin x cos x = 1 - 2 cos^2 x sin 2 x = sin^2 x

OpenStudy (mathmale):

Yes, that's what I expected. You've eliminated cos x, which is good. See any way of simplifying further?

OpenStudy (korosh23):

we should not have done that 2 sinx cos x = sin ^2 x sinx on both side could cancel out 2 = sin x / cos x 2 = tanx

OpenStudy (mathmale):

I was thrown a bit by your sin^2x. I read that as "the sine of 2 times x." Best to write "the square of sin x" as (sin x)^2 or \[\sin^2x\]

OpenStudy (korosh23):

yes sorry

OpenStudy (korosh23):

so am I right ?

OpenStudy (mathmale):

Rather than check your work further, I'll suggest you look for solutions of tan x = 2.

OpenStudy (mathmale):

Remember that tan x is defined for all x other than pi/2, pi+pi/2, and so on.

OpenStudy (mathmale):

So you may find several solutions within the interva [0, 2pi].

OpenStudy (mathmale):

Excuse me, [0, 2pi)

OpenStudy (korosh23):

yes tan x = 2 is positive in the first and third quadrant using the CAST rule so tan ^ -1 (2) = 1.107

OpenStudy (korosh23):

yes several solutions are present

OpenStudy (mathmale):

One way to get a handle on this is to graph tan x on the interval [0,2p) and then draw a horiz. line thru y=2. How many times does this line hit the curves of tan x?

OpenStudy (mathmale):

Recall that the period of tan x is pi (not 2pi).

OpenStudy (korosh23):

yes

OpenStudy (korosh23):

so it hit at 1.107 radian and 4.249 radian

OpenStudy (korosh23):

and 0

OpenStudy (mathmale):

It's important that you check your answers in the original equation. Set your calc to "RAD" mode and evaluate each side of the equ'n separately. Are the 2 sides =?

OpenStudy (korosh23):

Sorry I don't understand what you mean here

OpenStudy (mathmale):

It's important that you check your answers in the original equation. How would you do that? You've shared your roots in radian measure, so if you check them in the original equation, the calculator has to be set to RAD mode first.

OpenStudy (mathmale):

Are you Russian? I learned a semester's worth of Russian at UCLA in 1965. Seem to remember a word that sounds like khoroshoi, but can't remember its meaning.

OpenStudy (korosh23):

I see what you mean, yes my calculator has to be in radian mode at all times

OpenStudy (korosh23):

I am Persian :D

OpenStudy (korosh23):

I know there are common names in languages

OpenStudy (korosh23):

so there are 4 answers but I got 2

OpenStudy (korosh23):

these are the four answers     x = 0 , pi , 1.107, 4.249

OpenStudy (mathmale):

We might have to go back to the orignal equation and double-check the substitutions we made. It's promising that you've obtained two that are correct. I believe you cancelled sin x or cos x. Note that if you actually did that, you may have lost a source of additional solutions inadvertently. Persian, as in Iranian?

OpenStudy (korosh23):

yeah I canceled something out. I should not have done that. Yes Persian and Iranian are the same thing. Hey, you know that :D

OpenStudy (mathmale):

note that sin x=0 has roots 0 and pi within the interval in question, and that cos x=0 has the roots pi/2 and 3pi/2

OpenStudy (mathmale):

Are you currently in the USA, or in Iran?

OpenStudy (korosh23):

Canada

OpenStudy (mathmale):

Neat. Fervently wish that the US and Iran could improve their relationship.

OpenStudy (mathmale):

Could I possibly be of any further help with this particular trig problem?

OpenStudy (korosh23):

Yeah, there I hope too

OpenStudy (korosh23):

well give me a sec I will tell you

OpenStudy (korosh23):

when we have 2 sin x cos x = 1- cos ^2 x 2 sinx cos x = sin ^2 x what should I do after here?

OpenStudy (mathmale):

Try factoring out sin x from both sides, but do not discard that factor. What's left?

OpenStudy (mathmale):

Or, rather, what are the resulting factors?

OpenStudy (korosh23):

sinx (2cosx) = sinx ( six)

OpenStudy (korosh23):

sin x (2 cos x ) = sin x (sin x)

OpenStudy (mathmale):

yes, or (sin x)(2cos x-sin x)=0 that's not too promising, though.

OpenStudy (korosh23):

I believe yes we got the answer

OpenStudy (korosh23):

2 sin x cos x = sin^2 x sin^2x - 2 sinx cos x= 0 sinx ( sinx - 2 cos x) = 0 sin x = 0 and sin x = 2 cos x which is 2 = tan x so sin ^ -1 (0 ) = 0 and pi?

OpenStudy (mathmale):

What you've done looks good, correct. "the angle whose sine is 0" is 0, but also pi. The inverse sine of x at x=0 is {0, pi}.

OpenStudy (korosh23):

Fantastic, nice we solved together. Sorry if it took long time. Thanks for backup :D

OpenStudy (mathmale):

Very happy to have this opp to work with you. Later, perhaps? Take care.

OpenStudy (korosh23):

sure we should solve more questions in the future, you too